may you help me solve 2x+3y=13 3x+5y=10
Answer by Maths68(1474)

2x+3y=13 .........(1)
3x+5y=10 ..........(2)
Multiply (1) by 3 and (2) by 2
6x+9y=39.........(3)
6x+10y=20........(4)
Subtract (4) from (3)
6x+9y=39
-6x-10y=-20
------------------
-y=19
Multiply above equation by (-1)
(-1)(-y)=(-1)(19)
y=-19
Put the value of y in (1)
2x+3y=13 .........(1)
2x+3(-19)=13
2x-57=13
2x=13+57
2x=70
2x/2=70/2
x=35
x,y=35, -19


Check
=====
Put the values of x and y in (1)
2x+3y=13 .........(1)
2(35)+3(-19)=13
70-57=13
13=13
Put the values of x and y in (2)
3x+5y=10 ..........(2)
3(35)+5(-19)=10
105-95=10
10=10